# Exact overlaps for ”all” integrable matrix product states of rational spin chains

### v1

@@ ACKNOWLEDGMENTS
We just saw that for certain Hilbert spaces, the extra selection rules lead to zero pseudo-vacuum overlap, i.e., $\mathbf{B}=\langle\Psi|0\rangle=0$. Our general overlap formula (S-1) is still valid, but the situation is complicated by the fact that the $\beta$-parameters are not eigenvalues of the operator $\mathbf{B}$. In these cases, the closed formula for the $\beta$-parameters is not (yet) known. Since these are simple numbers (not depending on the Bethe roots), their determination is not too complicated in the concrete cases. They can be determined by a simple numerical analysis.
@@
In the $\mathfrak{g}=\mathfrak{gl}_{N}$ case there exist chiral or achiral pair structure.
@@
$\mathbf{F}^{(k)}(u)=\left[\mathbf{G}^{(k)}(u)\right]^{-1}\mathbf{G}^{(k+1)}(u),$ (S-23)
@@
$\mathbf{F}^{(k)}(u)=\mathbf{A}\mathrm{diag}(\mathcal{F}_{1}^{(k)}(u),\dots,\mathcal{F}_{d_{B}}^{(k)}(u))\mathbf{A}^{-1},$ (S-24)
where $\mathbf{A}$ is an invertible matrix of the boundary space. The final $\tilde{\mathcal{F}}$ functions defined as
~$\tilde{\mathcal{F}}_{j}^{(k)}(u)=\mathcal{F}_{j}^{(k)}(-iu-k/2)\sqrt{\frac{u^{2}}{u^{2}+1/4}}.$ (S-25)
For achiral pair structure we introduce $n_{+}+1=\left\lfloor\frac{N}{2}\right\rfloor+1$ commuting operators $\mathbf{G}^{(k)}(u)$, for $k=1,\dots,n_{+}+1$. The $F$-operators and their spectrum are defined in the same way as before (S-23), (S-24) for $k=1,\dots,n_{+}$. The final $\tilde{\mathcal{F}}$ functions defined as
~$\tilde{\mathcal{F}}_{j}^{(k)}(u)=\begin{cases}\mathcal{F}_{j}^{(k)}(-iu+\frac{N-2k}{4}),&k<\frac{N-1}{2}\\ \mathcal{F}_{j}^{(\frac{N-1}{2})}(-iu+\frac{1}{4})\sqrt{\frac{u+i/4}{u-i/4}},&k=\frac{N-1}{2}\\ \mathcal{F}_{j}^{(\frac{N}{2})}(-iu)\sqrt{\frac{u^{2}}{u^{2}+1/4}},&k=\frac{N}{2}\end{cases}$ (S-26)
@@
$\mathbf{B}=\mathbf{A}\mathrm{diag}(\beta_{1},\dots,\beta_{d_{B}})\mathbf{A}^{-1}.$ (S-27)
@@
For the twisted Yangian $Y^{+}(N)\equiv Y(\mathfrak{gl}_{N},\mathfrak{o}_{N})$ there is no extra selection rule. We define a series of nested K-matrices with the following recursion
$\mathbf{K}_{a,b}^{(k+1)}(u)=\mathbf{K}_{a,b}^{(k)}(u)-\mathbf{K}_{a,k}^{(k)}(u)\left[\mathbf{K}_{k,k}^{(k)}(u)\right]^{-1}\mathbf{K}_{k,b}^{(k)}(u),$ (S-28)
@@
S-III.2 $Y(\mathfrak{gl}_{2n},\mathfrak{sp}_{2n})$ (chiral)
S-III.2.1 $n=1$
We start with the $n=1$ case, i.e. $Y^{-}(2)\equiv Y(\mathfrak{gl}_{2},\mathfrak{sp}_{2})$. At first we have to find the factorization form
$\mathbf{K}_{a,b}(u)=\sum_{c,d=1}^{2}\mathbf{L}_{c,a}(u)\epsilon_{c,d}\mathbf{L}_{d,b}(-u)$ (S-30)
@@
$\tilde{\mathbf{K}}_{a,b}(u)=\mathbf{K}_{a,b}(u)+(u+1/2)\mathbf{L}_{1,a}(u)\mathbf{L}_{1,b}(-u),$ (S-31)
where $\mathfrak{b}\in\mathbb{C}$ is parameter. Now we can define the G-operators
$\begin{split}\mathbf{G}^{(1)}(u)&=\tilde{\mathbf{K}}_{1,1}(u),\\ \mathbf{G}^{(2)}(u)&=\tilde{\mathbf{K}}_{2,2}(u)-\tilde{\mathbf{K}}_{2,1}(u)\left[\tilde{\mathbf{K}}_{1,1}(u)\right]^{-1}\tilde{\mathbf{K}}_{1,2}(u).\end{split}$ (S-32)
S-III.2.2 General $n$
For general $n$, i.e. $Y^{-}(2n)\equiv Y(\mathfrak{gl}_{2n},\mathfrak{sp}_{2n})$, the nesting is defined as
$\mathbf{K}_{a,b}^{(k+1)}(u)=\mathbf{K}_{a,b}^{(k)}(u)-\sum_{\alpha,\beta=2k-1}^{2k}\mathbf{K}_{a,\alpha}^{(k)}(u)\widehat{\mathbf{K}}_{\alpha,\beta}^{(k)}(u)\mathbf{K}_{\beta,b}^{(k)}(u),$ (S-33)
for $a,b=2k+1,\dots,2n$. We defined the inverse matrix $\widehat{\mathbf{K}}_{\alpha,\beta}^{(k)}(z)$ for $\alpha,\beta=2k-1,2k$ as
$\sum_{\beta=2k-1}^{2k}\widehat{\mathbf{K}}_{\alpha,\beta}^{(k)}(z)\mathbf{K}_{\beta,\gamma}^{(k)}(z)=\delta_{\alpha,\gamma}.$ (S-34)
We can select $n$ commuting $Y^{-}(2)$ subalgebras
$\mathbf{k}^{(s)}(u)=\left(\begin{array}[]{cc}\mathbf{K}_{2s-1,2s-1}^{(s)}(u)&\mathbf{K}_{2s-1,2s}^{(s)}(u)\\ \mathbf{K}_{2s,2s-1}^{(s)}(u)&\mathbf{K}_{2s,2s}^{(s)}(u)\end{array}\right),$
where $s=1,\dots,n$. We can repeat the $Y^{-}(2)$ method separately. We have to find the factorization form
$\mathbf{k}_{a,b}^{(s)}(u)=\sum_{c,d=1}^{2}\mathbf{L}_{c,a}(u+s-1)\epsilon_{c,d}\mathbf{L}_{d,b}(-u-s+1)$ (S-35)
@@
$\tilde{\mathbf{k}}_{a,b}^{(s)}(u)=\mathbf{k}_{a,b}^{(s)}(u)+(u+s-1/2)\mathbf{L}_{1,a}(u+s-1)\mathbf{L}_{1,b}(-u-s+1)$ (S-36)
@@
$\begin{split}\mathbf{G}^{(2s-1)}(u)&=\tilde{\mathbf{k}}_{1,1}^{(s)}(u),\\ \mathbf{G}^{(2s)}(z)&=\tilde{\mathbf{k}}_{2,2}^{(s)}(u)-\tilde{\mathbf{k}}_{2,1}^{(s)}(u)\left[\tilde{\mathbf{k}}_{1,1}^{(s)}(u)\right]^{-1}\tilde{\mathbf{k}}_{1,2}^{(s)}(u),\end{split}$ (S-37)
for $s=1,\dots,n$.
@@
$\mathbf{K}=\left(\begin{array}[]{ccccccc}\mathbf{K}_{-n,-n}&\dots&\mathbf{K}_{-n,-1}&\mathbf{K}_{-n,0}&\mathbf{K}_{-n,1}&\dots&\mathbf{K}_{-n,n}\\ \vdots&\ddots&\vdots&\vdots&\vdots&\iddots&\vdots\\ \mathbf{K}_{-1,-n}&\dots&\mathbf{K}_{-1,-1}&\mathbf{K}_{-1,0}&\mathbf{K}_{-1,1}&\dots&\mathbf{K}_{-1,n}\\ \mathbf{K}_{0,-n}&\dots&\mathbf{K}_{0,-1}&\mathbf{K}_{0,0}&\mathbf{K}_{0,1}&\dots&\mathbf{K}_{0,n}\\ \mathbf{K}_{1,-n}&\dots&\mathbf{K}_{1,-1}&\mathbf{K}_{1,0}&\mathbf{K}_{1,1}&\dots&\mathbf{K}_{1,n}\\ \vdots&\iddots&\vdots&\vdots&\vdots&\ddots&\vdots\\ \mathbf{K}_{n,-n}&\dots&\mathbf{K}_{n,-1}&\mathbf{K}_{n,0}&\mathbf{K}_{n,1}&\dots&\mathbf{K}_{n,n}\end{array}\right)$ (S-38)
@@
$\mathbf{K}=\left(\begin{array}[]{cccccc}\mathbf{K}_{-n,-n}&\dots&\mathbf{K}_{-n,-1}&\mathbf{K}_{-n,1}&\dots&\mathbf{K}_{-n,n}\\ \vdots&\ddots&\vdots&\vdots&\iddots&\vdots\\ \mathbf{K}_{-1,-n}&\dots&\mathbf{K}_{-1,-1}&\mathbf{K}_{-1,1}&\dots&\mathbf{K}_{-1,n}\\ \mathbf{K}_{1,-n}&\dots&\mathbf{K}_{1,-1}&\mathbf{K}_{1,1}&\dots&\mathbf{K}_{1,n}\\ \vdots&\iddots&\vdots&\vdots&\ddots&\vdots\\ \mathbf{K}_{n,-n}&\dots&\mathbf{K}_{n,-1}&\mathbf{K}_{n,1}&\dots&\mathbf{K}_{n,n}\end{array}\right)$ (S-39)
@@
When $M=\left\lfloor\frac{N}{2}\right\rfloor$ there is no extra selection rule. We have a series of $K$-matrices $\mathbf{K}^{(k)}$ for $k=1,\dots,n$ with the recursion
@@
where $a,b=-n+s,-n+s+1,\dots,n-s-1,n-s$, with the initial condition $\mathbf{K}^{(1)}(u)\equiv\mathbf{K}(u)$. We also define the G-operators $\mathbf{G}^{(k)}$ for $k=1,\dots,n$ with
$\mathbf{G}^{(k)}(u)=\mathbf{K}_{n-k+1,-n+k-1}^{(k)}(u).$ (S-41)
@@
$\mathbf{G}^{(n+1)}(u)=\begin{cases}\mathbf{K}_{0,0}^{(n)}(u)-\mathbf{K}_{0,-1}^{(k)}(u)\left[\mathbf{K}_{1,-1}^{(n)}(u)\right]^{-1}\mathbf{K}_{1,0}^{(k)}(u)\Biggr{|}_{\mathfrak{b}_{n}=1},&\text{for }N=2n+1,\\ \mathbf{K}_{-1,1}^{(n)}(u)-\mathbf{K}_{-1,-1}^{(k)}(u)\left[\mathbf{K}_{1,-1}^{(n)}(u)\right]^{-1}\mathbf{K}_{1,1}^{(k)}(u)\Biggr{|}_{\mathfrak{b}_{n}=1},&\text{for }N=2n,\end{cases}$ (S-42)
@@
At first we have to find the factorization form
$\mathbf{K}(u)=\mathbf{L}^{-1}(u)\mathbf{L}(-u)$ (S-43)
@@
$\tilde{\mathbf{K}}_{a,b}(u)=\mathbf{K}_{a,b}(u)+u\sum_{c=1}^{n}\left[\mathbf{L}^{-1}(u)\right]_{a,c}\mathbf{L}_{-c,b}(-u)$ (S-44)
Now we can use the nesting (S-40)with the initial condition $\mathbf{K}^{(1)}(u)\equiv\tilde{\mathbf{K}}(u)$ to obtain the nested K-matrices $\mathbf{K}^{(k)}$ for $k=1,\dots,n$. We define the $G$-operators with (S-41), (S-42).
@@
For general $M$ we combine the methods of sections S-III.3.1 and S-III.3.2. We have a series of $K$-matrices $\mathbf{K}^{(k)}$ for $k=1,\dots,M+1$ with the recursion (S-40) and initial condition $\mathbf{K}^{(1)}(u)\equiv\mathbf{K}(u)$. We also define the G-operators $\mathbf{G}^{(k)}$ for $k=1,\dots,M$ with (S-41).
@@
$\tilde{\mathbf{K}}_{a,b}^{(M+1)}(u)=\mathbf{K}_{a,b}^{(M+1)}(u)+u\sum_{c=1}^{n-M}\left[\mathbf{L}^{-1}(u)\right]_{a,c}\mathbf{L}_{-c,b}(-u)$ (S-46)
We can continue the series of $K$-matrices $\mathbf{K}^{(k)}$ for $k=M+2,\dots,n$ with the recursion (S-40) and initial condition $\mathbf{K}^{(M+1)}(u)\equiv\tilde{\mathbf{K}}^{(M+1)}$. We define the remaining $G$-operators as $\mathbf{G}^{(M+1)}(u)=\tilde{\mathbf{K}}_{n-M,-n+M}^{(M+1)}(u)$ and with the equation (S-41) for $k=M+2,M+3,\dots,n$ and the $\mathbf{G}^{(n+1)}(u)$ with (S-42).
@@
~$\tilde{\mathcal{F}}_{j}^{(k)}(u)=\begin{cases}\mathcal{F}_{j}^{(k)}(-iu+\frac{n-k-1}{2})\sqrt{\frac{u^{2}}{u^{2}+1/4}},&k=1,\dots,n-2,\\ \mathcal{F}_{j}^{(k)}(-iu)\sqrt{\frac{u^{2}}{u^{2}+1/4}}&k=n-1,n\end{cases}$ (S-48)
For achiral pair structure we introduce $n$ commuting operators $\mathbf{G}^{(k)}(u)$ for $k=1,\dots,n$. The $F$-operators are defined as (S-23) for $k=1,\dots,n-1$. The final $\tilde{\mathcal{F}}$ functions defined as
~$\tilde{\mathcal{F}}_{j}^{(k)}(u)=\begin{cases}\mathcal{F}_{j}^{(k)}(-iu+\frac{n-k-1}{2})\sqrt{\frac{u^{2}}{u^{2}+1/4}},&k=1,\dots,n-2,\\ \mathcal{F}_{j}^{(n-1)}(-iu)\frac{u}{u-i/2},&k=n-1.\end{cases}$ (S-49)
@@
$\sum_{\beta=n-2s+1}^{n-2s+2}\widehat{\mathbf{K}}_{-\alpha,\beta}^{(s)}(z)\mathbf{K}_{\beta,-\gamma}^{(s)}(z)=\delta_{\alpha,\gamma}.$ (S-51)
@@
$\mathbf{k}^{(s)}(u)=\left(\begin{array}[]{cc}\mathbf{K}_{n-2s+2,-n+2s-2}^{(s)}(u)&\mathbf{K}_{n-2s+2,-n+2s-1}^{(s)}(u)\\ \mathbf{K}_{n-2s+1,-n+2s-2}^{(s)}(u)&\mathbf{K}_{n-2s+1,-n+2s-1}^{(s)}(u)\end{array}\right),$
@@
$\begin{split}\mathbf{k}_{a,b}^{(s)}(u)&=\sum_{c,d=1}^{2}\mathbf{L}_{c,a}(u-\kappa_{N-4s+4})\epsilon_{c,d}\mathbf{L}_{d,b}(-u),\\ \tilde{\mathbf{k}}_{a,b}^{(s)}(u)&=\mathbf{k}_{a,b}^{(s)}(u)+(u+1/2-\kappa_{N-4s+4}/2)\mathbf{L}_{1,a}(u-\kappa_{N-4s+4})\mathbf{L}_{1,b}(-u),\end{split}$ (S-52)
@@
$\mathbf{F}^{(s)}(u)=\left[\mathbf{G}^{(s)}(u)\right]^{-1}\mathbf{G}^{(s+1)}(u),$ (S-53)

### v2

@@ Acknowledgments
We just saw that for certain Hilbert spaces, the extra selection rules lead to zero pseudo-vacuum overlap, i.e., $\mathbf{B}=\langle\Psi|0\rangle=0$. Our general overlap formula (S-1) is still valid, but for these Hilbert spaces the $\beta$-parameters are not eigenvalues of the operator $\mathbf{B}$. In these cases, the closed formula for the $\beta$-parameters is not known (yet). Since these are simple numbers (do not depending on the Bethe roots), their determination is not too complicated in the concrete cases. They can be determined by a simple numerical analysis.
@@
In the $\mathfrak{g}=\mathfrak{gl}_{N}$ case there exist chiral or achiral pair structures.
@@
$\mathbf{F}^{(k)}(u)=\left[\mathbf{G}^{(k)}(u)\right]^{-1}\mathbf{G}^{(k+1)}(u),$ (S-22)
@@
$\mathbf{F}^{(k)}(u)=\mathbf{A}\mathrm{diag}(\mathcal{F}_{1}^{(k)}(u),\dots,\mathcal{F}_{d_{B}}^{(k)}(u))\mathbf{A}^{-1},$ (S-23)
where $\mathbf{A}$ is an invertible matrix of the boundary space. The final $\tilde{\mathcal{F}}$ functions are defined as
~$\tilde{\mathcal{F}}_{j}^{(k)}(u)=\mathcal{F}_{j}^{(k)}(iu-k/2)\sqrt{\frac{u^{2}}{u^{2}+1/4}}.$ (S-24)
For achiral pair structure we introduce $n_{+}+1=\left\lfloor\frac{N}{2}\right\rfloor+1$ commuting operators $\mathbf{G}^{(k)}(u)$, for $k=1,\dots,n_{+}+1$. The $F$-operators and their spectrum are defined in the same way as before (S-22), (S-23) for $k=1,\dots,n_{+}$. The final $\tilde{\mathcal{F}}$ functions defined as
~$\tilde{\mathcal{F}}_{j}^{(k)}(u)=\begin{cases}\mathcal{F}_{j}^{(k)}(iu+\frac{N-2k}{4}),&k<\frac{N-1}{2},\\ \mathcal{F}_{j}^{(\frac{N-1}{2})}(iu+\frac{1}{4})\sqrt{\frac{u-i/4}{u+i/4}},&k=\frac{N-1}{2},\\ \mathcal{F}_{j}^{(\frac{N}{2})}(iu)\sqrt{\frac{u^{2}}{u^{2}+1/4}},&k=\frac{N}{2}.\end{cases}$ (S-25)
@@
$\mathbf{B}=\mathbf{A}\mathrm{diag}(\beta_{1},\dots,\beta_{d_{B}})\mathbf{A}^{-1}.$ (S-26)
@@
For the twisted Yangian $Y^{+}(N)\equiv Y(\mathfrak{gl}_{N},\mathfrak{o}_{N})$ we define a series of nested K-matrices with the following recursion
$\mathbf{K}_{a,b}^{(k+1)}(u)=\mathbf{K}_{a,b}^{(k)}(u)-\mathbf{K}_{a,k}^{(k)}(u)\left[\mathbf{K}_{k,k}^{(k)}(u)\right]^{-1}\mathbf{K}_{k,b}^{(k)}(u),$ (S-27)
@@
S-III.2 $Y(\mathfrak{gl}_{2k},\mathfrak{sp}_{2k})$ (chiral)
S-III.2.1 $k=1$
We start with the $k=1$ case, i.e. $Y^{-}(2)\equiv Y(\mathfrak{gl}_{2},\mathfrak{sp}_{2})$. At first we have to find the factorization form
$\mathbf{K}_{a,b}(u)=\sum_{c,d=1}^{2}\mathbf{L}_{c,a}(u)\epsilon_{c,d}\mathbf{L}_{d,b}(-u).$ (S-29)
@@
$\tilde{\mathbf{K}}_{a,b}(u)=\mathbf{K}_{a,b}(u)+(u+1/2)\mathbf{L}_{1,a}(u)\mathbf{L}_{1,b}(-u).$ (S-30)
Now we can define the G-operators
$\begin{split}\mathbf{G}^{(1)}(u)&=\tilde{\mathbf{K}}_{1,1}(u),\\ \mathbf{G}^{(2)}(u)&=\tilde{\mathbf{K}}_{2,2}(u)-\tilde{\mathbf{K}}_{2,1}(u)\left[\tilde{\mathbf{K}}_{1,1}(u)\right]^{-1}\tilde{\mathbf{K}}_{1,2}(u).\end{split}$ (S-31)
S-III.2.2 General $k$
For general $k$, i.e. $Y^{-}(2k)\equiv Y(\mathfrak{gl}_{2k},\mathfrak{sp}_{2k})$, the nesting is defined as
$\mathbf{K}_{a,b}^{(s+1)}(u)=\mathbf{K}_{a,b}^{(s)}(u)-\sum_{\alpha,\beta=2s-1}^{2s}\mathbf{K}_{a,\alpha}^{(s)}(u)\widehat{\mathbf{K}}_{\alpha,\beta}^{(s)}(u)\mathbf{K}_{\beta,b}^{(s)}(u),$ (S-32)
for $a,b=2s+1,\dots,2k$. We defined the inverse matrix $\widehat{\mathbf{K}}_{\alpha,\beta}^{(s)}(u)$ for $\alpha,\beta=2s-1,2s$ as
$\sum_{\beta=2s-1}^{2s}\widehat{\mathbf{K}}_{\alpha,\beta}^{(s)}(u)\mathbf{K}_{\beta,\gamma}^{(s)}(u)=\delta_{\alpha,\gamma}.$ (S-33)
We can select $k$ commuting $Y^{-}(2)$ subalgebras
$\mathbf{k}^{(s)}(u)=\left(\begin{array}[]{cc}\mathbf{K}_{2s-1,2s-1}^{(s)}(u)&\mathbf{K}_{2s-1,2s}^{(s)}(u)\\ \mathbf{K}_{2s,2s-1}^{(s)}(u)&\mathbf{K}_{2s,2s}^{(s)}(u)\end{array}\right),$ (S-34)
where $s=1,\dots,k$. We can repeat the $Y^{-}(2)$ method separately. We have to find the factorization form
$\mathbf{k}_{a,b}^{(s)}(u)=\sum_{c,d=1}^{2}\mathbf{L}_{c,a}(u+s-1)\epsilon_{c,d}\mathbf{L}_{d,b}(-u-s+1),$ (S-35)
@@
$\tilde{\mathbf{k}}_{a,b}^{(s)}(u)=\mathbf{k}_{a,b}^{(s)}(u)+(u+s-1/2)\mathbf{L}_{1,a}(u+s-1)\mathbf{L}_{1,b}(-u-s+1).$ (S-36)
@@
$\begin{split}\mathbf{G}^{(2s-1)}(u)&=\tilde{\mathbf{k}}_{1,1}^{(s)}(u),\\ \mathbf{G}^{(2s)}(u)&=\tilde{\mathbf{k}}_{2,2}^{(s)}(u)-\tilde{\mathbf{k}}_{2,1}^{(s)}(u)\left[\tilde{\mathbf{k}}_{1,1}^{(s)}(u)\right]^{-1}\tilde{\mathbf{k}}_{1,2}^{(s)}(u),\end{split}$ (S-37)
for $s=1,\dots,k$.
@@
$\mathbf{K}=\left(\begin{array}[]{ccccccc}\mathbf{K}_{-n,-n}&\dots&\mathbf{K}_{-n,-1}&\mathbf{K}_{-n,0}&\mathbf{K}_{-n,1}&\dots&\mathbf{K}_{-n,n}\\ \vdots&\ddots&\vdots&\vdots&\vdots&\iddots&\vdots\\ \mathbf{K}_{-1,-n}&\dots&\mathbf{K}_{-1,-1}&\mathbf{K}_{-1,0}&\mathbf{K}_{-1,1}&\dots&\mathbf{K}_{-1,n}\\ \mathbf{K}_{0,-n}&\dots&\mathbf{K}_{0,-1}&\mathbf{K}_{0,0}&\mathbf{K}_{0,1}&\dots&\mathbf{K}_{0,n}\\ \mathbf{K}_{1,-n}&\dots&\mathbf{K}_{1,-1}&\mathbf{K}_{1,0}&\mathbf{K}_{1,1}&\dots&\mathbf{K}_{1,n}\\ \vdots&\iddots&\vdots&\vdots&\vdots&\ddots&\vdots\\ \mathbf{K}_{n,-n}&\dots&\mathbf{K}_{n,-1}&\mathbf{K}_{n,0}&\mathbf{K}_{n,1}&\dots&\mathbf{K}_{n,n}\end{array}\right).$ (S-38)
@@
$\mathbf{K}=\left(\begin{array}[]{cccccc}\mathbf{K}_{-n,-n}&\dots&\mathbf{K}_{-n,-1}&\mathbf{K}_{-n,1}&\dots&\mathbf{K}_{-n,n}\\ \vdots&\ddots&\vdots&\vdots&\iddots&\vdots\\ \mathbf{K}_{-1,-n}&\dots&\mathbf{K}_{-1,-1}&\mathbf{K}_{-1,1}&\dots&\mathbf{K}_{-1,n}\\ \mathbf{K}_{1,-n}&\dots&\mathbf{K}_{1,-1}&\mathbf{K}_{1,1}&\dots&\mathbf{K}_{1,n}\\ \vdots&\iddots&\vdots&\vdots&\ddots&\vdots\\ \mathbf{K}_{n,-n}&\dots&\mathbf{K}_{n,-1}&\mathbf{K}_{n,1}&\dots&\mathbf{K}_{n,n}\end{array}\right).$ (S-39)
@@
When $M=\left\lfloor\frac{N}{2}\right\rfloor$ we have a series of $K$-matrices $\mathbf{K}^{(s)}$ for $s=1,\dots,n$ with the recursion
@@
where $a,b=-n+s,-n+s+1,\dots,n-s-1,n-s$, with the initial condition $\mathbf{K}^{(1)}(u)\equiv\mathbf{K}(u)$. We also define the G-operators $\mathbf{G}^{(s)}$ for $s=1,\dots,n$ with
$\mathbf{G}^{(s)}(u)=\mathbf{K}_{n-s+1,-n+s-1}^{(s)}(u),$ (S-41)
@@
$\mathbf{G}^{(n+1)}(u)=\begin{cases}\mathbf{K}_{0,0}^{(n)}(u)-\mathbf{K}_{0,-1}^{(n)}(u)\left[\mathbf{K}_{1,-1}^{(n)}(u)\right]^{-1}\mathbf{K}_{1,0}^{(n)}(u),&\text{for }N=2n+1,\\ \mathbf{K}_{-1,1}^{(n)}(u)-\mathbf{K}_{-1,-1}^{(n)}(u)\left[\mathbf{K}_{1,-1}^{(n)}(u)\right]^{-1}\mathbf{K}_{1,1}^{(n)}(u),&\text{for }N=2n.\end{cases}$ (S-42)
@@
For $M=0$. at first we have to find the factorization form
$\mathbf{K}(u)=\mathbf{L}^{-1}(u)\mathbf{L}(-u).$ (S-43)
@@
$\tilde{\mathbf{K}}_{a,b}(u)=\mathbf{K}_{a,b}(u)+u\sum_{c=1}^{n}\left[\mathbf{L}^{-1}(u)\right]_{a,c}\mathbf{L}_{-c,b}(-u).$ (S-44)
Now we can use the nesting (S-40) with the initial condition $\mathbf{K}^{(1)}(u)\equiv\tilde{\mathbf{K}}(u)$ to obtain the nested K-matrices $\mathbf{K}^{(s)}$ for $s=1,\dots,n$. We define the $G$-operators with (S-41), (S-42).
@@
For general $M$ we combine the methods of sections S-III.3.1 and S-III.3.2. We have a series of $K$-matrices $\mathbf{K}^{(s)}$ for $s=1,\dots,M+1$ with the recursion (S-40) and initial condition $\mathbf{K}^{(1)}(u)\equiv\mathbf{K}(u)$. We also define the G-operators $\mathbf{G}^{(s)}$ for $s=1,\dots,M$ with (S-41).
@@
$\tilde{\mathbf{K}}_{a,b}^{(M+1)}(u)=\mathbf{K}_{a,b}^{(M+1)}(u)+u\sum_{c=1}^{n-M}\left[\mathbf{L}^{-1}(u)\right]_{a,c}\mathbf{L}_{-c,b}(-u).$ (S-46)
We can continue the series of $K$-matrices $\mathbf{K}^{(s)}$ for $s=M+2,\dots,n$ with the recursion (S-40) and initial condition $\mathbf{K}^{(M+1)}(u)\equiv\tilde{\mathbf{K}}^{(M+1)}$. We define the remaining $G$-operators as $\mathbf{G}^{(M+1)}(u)=\tilde{\mathbf{K}}_{n-M,-n+M}^{(M+1)}(u)$ and with the equation (S-41) for $s=M+2,M+3,\dots,n$ and the $\mathbf{G}^{(n+1)}(u)$ with (S-42).
@@
~$\tilde{\mathcal{F}}_{j}^{(k)}(u)=\begin{cases}\mathcal{F}_{j}^{(k)}(iu+\frac{n-k-1}{2})\sqrt{\frac{u^{2}}{u^{2}+1/4}},&k=1,\dots,n-2,\\ \mathcal{F}_{j}^{(k)}(iu)\sqrt{\frac{u^{2}}{u^{2}+1/4}},&k=n-1,n.\end{cases}$ (S-48)
For achiral pair structure we introduce $n$ commuting operators $\mathbf{G}^{(k)}(u)$ for $k=1,\dots,n$. The $F$-operators are defined as (S-22) for $k=1,\dots,n-1$. The final $\tilde{\mathcal{F}}$ functions defined as
~$\tilde{\mathcal{F}}_{j}^{(k)}(u)=\begin{cases}\mathcal{F}_{j}^{(k)}(iu+\frac{n-k-1}{2})\sqrt{\frac{u^{2}}{u^{2}+1/4}},&k=1,\dots,n-2,\\ \mathcal{F}_{j}^{(n-1)}(iu)\frac{u}{u+i/2},&k=n-1.\end{cases}$ (S-49)
@@
$\sum_{\beta=n-2s+1}^{n-2s+2}\widehat{\mathbf{K}}_{-\alpha,\beta}^{(s)}(u)\mathbf{K}_{\beta,-\gamma}^{(s)}(u)=\delta_{\alpha,\gamma},$ (S-51)
@@
$\mathbf{k}^{(s)}(u)=\left(\begin{array}[]{cc}\mathbf{K}_{n-2s+2,-n+2s-2}^{(s)}(u)&\mathbf{K}_{n-2s+2,-n+2s-1}^{(s)}(u)\\ \mathbf{K}_{n-2s+1,-n+2s-2}^{(s)}(u)&\mathbf{K}_{n-2s+1,-n+2s-1}^{(s)}(u)\end{array}\right),$ (S-52)
@@
$\begin{split}\mathbf{k}_{a,b}^{(s)}(u)&=\sum_{c,d=1}^{2}\mathbf{L}_{c,a}(u-\kappa_{N-4s+4})\epsilon_{c,d}\mathbf{L}_{d,b}(-u),\\ \tilde{\mathbf{k}}_{a,b}^{(s)}(u)&=\mathbf{k}_{a,b}^{(s)}(u)+(u+1/2-\kappa_{N-4s+4}/2)\mathbf{L}_{1,a}(u-\kappa_{N-4s+4})\mathbf{L}_{1,b}(-u),\end{split}$ (S-53)
@@
$\mathbf{F}^{(s)}(u)=\left[\mathbf{G}^{(s)}(u)\right]^{-1}\mathbf{G}^{(s+1)}(u),$ (S-54)